# When Trackers Date Fish: A Benchmark and Framework for Underwater Multiple Fish Tracking

### v1

@@ Abstract
Multiple object tracking (MOT) technology has made significant progress in terrestrial applications, but underwater tracking scenarios remain underexplored despite their importance to marine ecology and aquaculture. We present Multiple Fish Tracking Dataset 2025 (MFT25), the first comprehensive dataset specifically designed for underwater multiple fish tracking, featuring 15 diverse video sequences with 408,578 meticulously annotated bounding boxes across 48,066 frames. Our dataset captures various underwater environments, fish species, and challenging conditions including occlusions, similar appearances, and erratic motion patterns. Additionally, we introduce Scale-aware and Unscented Tracker (SU-T), a specialized tracking framework featuring an Unscented Kalman Filter (UKF) optimized for non-linear fish swimming patterns and a novel Fish-Intersection-over-Union (FishIoU) matching that accounts for the unique morphological characteristics of aquatic species. Extensive experiments demonstrate that our SU-T baseline achieves state-of-the-art performance on MFT25, with 34.1 HOTA and 44.6 IDF1, while revealing fundamental differences between fish tracking and terrestrial object tracking scenarios. MFT25 establishes a robust foundation for advancing research in underwater tracking systems with important applications in marine biology, aquaculture monitoring, and ecological conservation. The dataset and codes are released at https://vranlee.github.io/SU-T/.
Figure 1: Challenges, features, and performance of the MFT25 dataset and SU-T baseline.
@@ 1 Introduction
Current methods largely rely on data-driven approaches, leveraging high-precision detectors to obtaizn real-time target positions [48, 21]. However, tracking fish in complex underwater environments presents several challenges, as shown in Fig. 1(a). On the one hand, high morphological similarity among individual fish combined with their erratic movement patterns frequently leads to identity switches and trajectory fragmentation [38, 22]. On the other hand, existing public datasets suffer from insufficient diversity and poor image quality [35, 34], limiting the development of tracking models with strong generalization capabilities across complex scenarios.
To address these challenges, we present Multiple Fish Tracking Dataset 2025 (MFT25), a large-scale dataset specifically designed for underwater multiple fish object tracking, alongside Scale-aware and Unscented Tracker (SU-T), an efficient, lightweight baseline model for online tracking. Our dataset and tracker aim to establish a robust foundation for advancing research in underwater object tracking systems with practical applications in marine ecology and aquaculture. Our main contributions are summarized as follows:
  • We introduce MFT25, the first comprehensive multiple fish tracking dataset featuring 15 diverse video sequences with 408,578 meticulously annotated bounding boxes across 48,066 frames, capturing various underwater environments, fish species, and challenging conditions including occlusions, rapid direction changes, and visually similar appearances.
@@
  • We conduct extensive comparative experiments demonstrating that our tracker achieves state-of-the-art performance on MFT25, with 34.1 HOTA and 44.6 IDF1. Through quantitative analysis, we highlight the fundamental differences between fish tracking and land-based object tracking scenarios, as illustrated in Fig. 1(b) and (c).
@@ 2.1 Toward Fish Tracking
2.1.1 Related Methods
Fish tracking confronts distinctive challenges compared to terrestrial object tracking due to complex underwater environments, unique fish morphology, and erratic swimming behaviors [8]. Early approaches relied primarily on traditional computer vision techniques such as background subtraction [38] and object segmentation [17], which performed adequately only in controlled environments. Recent advances have shifted toward sophisticated deep learning implementations, including SiamRPN architectures [41], appearance-based models [23], graph-based tracking approaches [18], and advanced Swin Transformer architectures [47].
While MOT methods for terrestrial and underwater targets share fundamental principles, terrestrial tracking has been significantly more extensively developed [15]. Recent terrestrial MOT research has explored diverse technical aspects including multi-modal fusion strategies [20], adaptive frame rate processing [26, 12], computational efficiency optimizations [24], robust appearance modeling [37], depth information integration [27], and high-density scenario handling [19]. Contemporary MOT frameworks can be systematically categorized into three principal architectural approaches: Joint Detection and Embedding-based (JDE-based), Transformer-based, and Separate Detection and Embedding-based (SDE-based) methods.
JDE-based methods employ unified network architectures that simultaneously learn detection and tracking features, achieving training efficiency but often compromising feature specialization [42, 48, 21]. Transformer-based approaches leverage attention mechanisms to establish sophisticated spatiotemporal relationships, achieving high performance but introducing substantial computational overhead and memory requirements [11, 7, 44, 31, 22, 50]. SDE-based methods, predominantly extending the SORT framework [4], have gained widespread adoption by decoupling high-performance detectors from specialized appearance models, offering an optimal balance between accuracy and efficiency [49, 5, 43, 13, 1, 45].
Recent innovations in MOT have also explored camera pose estimation [46], generative diffusion models [29], and pixel-wise trajectory propagation [51]. However, these sophisticated techniques often demonstrate limited adaptability to underwater environments and insufficient real-time performance, rendering them suboptimal for practical fish tracking applications. The unique characteristics of underwater environments—including variable lighting conditions, water turbidity, and the distinctive motion patterns of aquatic species—necessitate specialized tracking solutions.
@@ 2.1.2 Related Datasets
While several fish-oriented video datasets exist, including Fish4Knowledge, DeepFish, and SeaCLEF [8], they present significant limitations for modern tracking applications. These early datasets typically suffer from low resolution, poor visibility conditions that obscure fish identities, and inconsistent annotation formats that impede effective model training. More recent standardized fish tracking datasets such as BrackishMOT [35], 3D-ZeF [34], and MFT22 [22] have emerged with consistent annotation protocols. However, these datasets remain limited in both environmental diversity and scale, typically featuring simplified scenarios with controlled conditions. The absence of comprehensive, high-quality, and standardized fish tracking datasets thus represents a critical bottleneck constraining significant advances in underwater MFT research.
@@ 2.2 Dataset Construction
To ensure comprehensive scenario coverage, we systematically deployed multiple camera configurations, including both overhead and horizontal perspectives, across varied illumination conditions from daylight to nocturnal settings. The dataset consists exclusively of authentic footage without synthetic augmentation, preserving the natural complexity of scenarios. All bounding box annotations were created using DarkLabel [^1] https://github.com/darkpgmr/DarkLabel software through manual selection and verification processes. The resulting MFT25 dataset encompasses 15 diverse video sequences containing 223 distinct fish trajectories across 48,066 frames, with a total of 408,578 precisely annotated bounding boxes. This represents a substantial advancement in scale, containing 2.6-9.1 times more annotated instances compared to previous fish tracking datasets. Table 1 presents a comprehensive statistical comparison with existing fish tracking benchmarks.
@@
Frames 12,466 28,800 9,100 48,066
Boxes 44,809 86,400 155,437 408,578
Moreover, we analyze the distinctive movement characteristics of fish in our dataset through detailed quantitative analysis in Section 4.4.
@@ 3.2.1 Sigma Points Generation
For state vector $\mathbf{x}\in\mathbb{R}^{n}$ with covariance $\mathbf{P}$, we generate $12n+$ 1 sigma points:
@@ 3.4 Assocation
$\mathcal{M}_{1},\mathcal{U}_{\mathcal{D}},\mathcal{U}_{\mathcal{T}}\leftarrow$ Hungarian($-\mathbf{C}$)
@@
for $(i,j)\in$ Hungarian($-\mathbf{C}_{last}$) where $\mathbf{C}_{last}[i,j]>\tau_{iou}$ do
@@
Remove tracks with $time\_since\_update>max\_age$
@@
4 Experiments
4.1 Implementation Details
For all experiments, we employed the same YOLOX-X [14] detector and the same Re-ID network [16] for SDE-based models, trained with consistent hyperparameter configurations, following the established protocols from ByteTrack [49] and BoT-SORT [1]. All models were trained on the MFT25 training set with a Tesla A100 GPU, with performance on the test set evaluated using standard MOT metrics, including the comprehensive HOTA [28] metric alongside traditional CLEAR [3] metrics such as MOTA, IDF1, and ID switches (IDs).
4.2 Benchmark Results
@@ 4.3 Ablation Studies
Table 6 presents the performance of our SU-T baseline across individual sequences in the MFT25 dataset, revealing substantial variation in tracking difficulty. The PF-002 sequence demonstrates exceptional performance with 65.2 HOTA, 98.7 IDF1, and none IDs, highlighting the effectiveness of SU-T under favorable conditions with distinct fish appearances and predictable movements. Similarly, BT-004 achieves strong results with 51.2 HOTA and 68.9 IDF1, demonstrating robust tracking capabilities in moderately challenging scenarios where fish maintain relatively consistent swimming patterns.
In contrast, sequences like SN-011 exhibit significantly lower performance with 20.3 HOTA, due to challenging conditions including numerous occlusions and highly erratic movements. MSK-003 shows moderate 34.2 HOTA but suffers from higher 134 IDs, indicating difficulties in maintaining consistent identities in crowded environments. SN-009 and SN-011 both demonstrate substantial identification failures, with high IDFN values of 30,853 and 23,575, respectively. This reveals the limitations of current tracking approaches in visibility-constrained underwater environments.
To visualize the practical effectiveness of SU-T, Fig. 5 presents tracking results across diverse scenarios in the MFT25 dataset. The visualizations demonstrate that SU-T maintains consistent identity assignments despite varying challenges including different fish densities, species morphologies, and environmental conditions.
@@
Clip HOTA↑ IDF1↑ MOTA↑ AssA↑ IDs↓ IDFP↓ IDFN↓
BT-002 35.961 58.562 55.923 28.788 35 1917 2095
BT-004 51.244 68.933 93.028 42.125 21 2299 2408
MSK-003 34.188 53.067 50.393 28.930 134 12488 15001
PF-002 65.162 98.650 97.300 65.399 0 82 80
SN-009 27.169 28.799 82.766 12.692 173 29339 30853
SN-011 20.264 26.165 47.423 10.421 161 27146 23575
SN-014 40.046 47.718 84.339 26.700 83 9840 10802
Total 33.351 41.717 68.450 22.425 607 83111 84814
@@
4.4 Statistical Analysis
We conducted comparative statistical analyses across diverse tracking datasets, including pedestrians [10], dancers [40], bees [6], and cells [2], to quantify the unique motion characteristics that distinguish fish tracking from other domains. Fig. 1(c) illustrates the distribution of motion directions, while Fig. 6 presents the angular velocity profiles across datasets. Our analysis reveals statistically significant differences in movement patterns. The MOT17 [10] dataset demonstrates remarkably stable angular velocities and predominantly linear motion patterns typical of pedestrian movement. In contrast, MFT25 exhibit substantially higher angular velocity variations, with more chaotic directional changes while maintaining consistently high swimming speeds. This erratic behavior stems from the hydrodynamic properties of underwater locomotion and environmental factors, creating inherently non-linear motion trajectories.
As shown in Fig. 7, MFT25 features comparatively higher average motion speeds, while the CTMC [2] dataset displays the lowest velocity profiles. These fundamental kinematic differences underscore why conventional tracking algorithms optimized for terrestrial scenarios often underperform in underwater environments, highlighting the need for specialized fish tracking approaches that can accommodate these unique movement characteristics.
5 Conclusion
In this paper, we introduces a unified underwater MFT benchmark and a specialized tracking framework for fish morphology and erratic swimming patterns. Our baseline achieves state-of-the-art performance with 34.1 HOTA, significantly outperforming other trackers. Statistical analysis reveals fundamental differences between fish and terrestrial tracking scenarios, highlighting the necessity for specialized underwater approaches. However, significant challenges remain in handling visually similar fish appearances, extreme density scenarios, and highly erratic swimming patterns, as evidenced by performance variations across sequences. Nevertheless, MFT25 and SU-T establish a robust foundation for advancing underwater tracking systems with important applications in marine biology, aquaculture monitoring, and ecological conservation.

### v3

@@ Abstract
Multiple object tracking (MOT) technology has made significant progress in terrestrial applications, but underwater tracking scenarios remain underexplored despite their importance to marine ecology and aquaculture. In this paper, we present Multiple Fish Tracking Dataset 2025 (MFT25), a comprehensive dataset specifically designed for underwater multiple fish tracking, featuring 15 diverse video sequences with 408,578 meticulously annotated bounding boxes across 48,066 frames. Our dataset captures various underwater environments, fish species, and challenging conditions including occlusions, similar appearances, and erratic motion patterns. Additionally, we introduce Scale-aware and Unscented Tracker (SU-T), a specialized tracking framework featuring an Unscented Kalman Filter (UKF) optimized for non-linear fish swimming patterns and a novel Fish-Intersection-over-Union (FishIoU) matching that accounts for the unique morphological characteristics of aquatic species. Extensive experiments demonstrate that our SU-T baseline achieves state-of-the-art performance on MFT25, with 34.1 HOTA and 44.6 IDF1, while revealing fundamental differences between fish tracking and terrestrial object tracking scenarios. The dataset and codes are released at https://vranlee.github.io/SU-T/.
Figure 1: Evaluation of various MFT and MOT methods on MFT25 benchmark. Detailed results are provided in Table 2.
@@ 1 Introduction
A tracker’s greatest challenge is not merely to find a fish, but to arrange a date with the same fleeting shadow—a perfect alignment of time and space.
@@
Figure 2: Distribution of target movement directions across datasets. Directional instability is notably more pronounced in the fish dataset compared to other target categories.
Current methods largely rely on data-driven approaches, leveraging high-precision detectors to obtain real-time target positions [51, 23]. However, tracking fish in complex underwater environments presents several challenges, as shown in Fig. 3. On the one hand, high morphological similarity among individual fish combined with their erratic movement patterns frequently leads to identity switches and trajectory fragmentation [40, 24]. On the other hand, existing public datasets suffer from insufficient diversity and poor image quality [37, 36], limiting the development of tracking models with strong generalization capabilities across complex scenarios.
Figure 3: The challenges of multiple fish tracking arise from factors such as fish physiological features and the complexity of underwater scenarios.
To address these challenges, we present Multiple Fish Tracking Dataset 2025 (MFT25), a large-scale dataset specifically designed for underwater MOT task, alongside Scale-aware and Unscented Tracker (SU-T), an efficient, lightweight baseline model for online tracking. Our dataset and tracker aim to establish a robust foundation for advancing research in underwater object tracking systems with practical applications in marine ecology and aquaculture. Our main contributions are summarized as follows:
  • We introduce MFT25, a large-scale fish dataset for MOT, featuring 15 diverse video sequences with 408,578 meticulously annotated bounding boxes across 48,066 frames, capturing various underwater environments, fish species, and challenging conditions including occlusions, rapid direction changes, and visually similar appearances.
@@
  • We conduct extensive comparative experiments demonstrating that our tracker achieves state-of-the-art performance on MFT25, with 34.1 HOTA and 44.6 IDF1, as illustrated in Fig. 1.
  • Through quantitative analysis, we highlight the fundamental differences between fish tracking and land-based object tracking scenarios, as shown in Fig. 2.
@@ 2.1 Toward Fish Tracking
Related Methods
Fish tracking presents unique challenges due to complex underwater environments, distinctive morphology, and erratic swimming behaviors [8]. Early approaches used traditional techniques like background subtraction [40] and object segmentation [18], while recent advances employ deep learning methods including SiamRPN [43], appearance-based models [25], graph-based tracking [19], and Swin Transformers [49]. While sharing principles with terrestrial tracking, underwater MOT remains less developed [16]. Terrestrial MOT research has explored multi-modal fusion [22], adaptive frame rates [28, 13], computational efficiency [26], appearance modeling [39], depth integration [29], and high-density scenarios [21].
Contemporary frameworks fall into three categories: JDE methods use unified networks achieving training efficiency but compromising specialization [44, 51, 23]; Transformer approaches leverage attention mechanisms with high performance but substantial overhead [12, 7, 46, 33, 24, 53]; SDE methods extend SORT [4], decoupling detectors from appearance models for balanced accuracy and efficiency [52, 5, 45, 14, 1, 47]. Recent innovations include camera pose estimation [48], generative diffusion models [31], and pixel-wise trajectory propagation [54], but show limited underwater adaptability and insufficient real-time performance for practical fish tracking applications.
@@
While several fish-oriented video datasets exist, including Fish4Knowledge, DeepFish, and SeaCLEF [8], they present significant limitations for modern tracking applications. These early datasets typically suffer from low resolution, poor visibility conditions that obscure fish identities, and inconsistent annotation formats that impede effective model training. Furthermore, other specialized datasets, such as FishTrack23 [9] and WebUOT-1M [50], are designed for Single Object Tracking (SOT), a fundamentally different task that provides annotations only for the initialized target. In contrast, MOT datasets require frame-by-frame annotations for all visible targets, along with consistent identity labels to enable long-term association. The CFC dataset [20] utilizes sonar imaging, which represents a distinct data modality from the optical videos commonly used in aquaculture applications. More recent standardized fish tracking datasets, including BrackishMOT [37], 3D-ZeF [36], and MFT22 [24], have emerged with consistent annotation protocols. However, these datasets remain limited in both environmental diversity and scale, typically featuring simplified scenarios under controlled conditions. The absence of comprehensive, high-quality, and standardized fish tracking datasets thus represents a critical bottleneck that constrains significant advances in underwater MFT research.
@@ 2.2 Dataset Construction
To ensure comprehensive scenario coverage, we systematically deployed multiple camera configurations, including both overhead and horizontal perspectives, across varied illumination conditions from daylight to nocturnal settings. The dataset consists exclusively of authentic footage without synthetic augmentation, preserving the natural complexity of scenarios. All bounding box annotations were created using DarkLabel [^1] https://github.com/darkpgmr/DarkLabel software through manual selection and verification processes. The resulting MFT25 dataset encompasses 15 diverse video sequences containing 223 distinct fish trajectories across 48,066 frames, with a total of 408,578 precisely annotated bounding boxes. This represents a substantial advancement in scale, containing 2.6-8.3 times more annotated instances compared to previous fish tracking datasets. Table 1 presents a comprehensive statistical comparison with existing fish tracking benchmarks.
@@
Frames 14,017 14,398 9,100 48,066
Boxes 49,364 86,452 155,437 408,578
@@ 3.1 Framework
3.2 Detector and Re-ID
Considering the variability of fish movements and the significant scale variations due to varying distances from the camera, our tracker adopts a mainstream pyramid-based design following [52, 5, 47] and employs decoupled heads to predict point centers, bounding boxes, offsets, and confidence scores [15]. The different decoupled heads share feature parameters across pyramid levels. Additionally, an optional re-identification module continuously updates GeM Pooling through learnable parameters, outputting target appearance features that are incorporated into the cost matrix for subsequent association, thereby providing robust appearance similarity cues for tracking.
@@ 3.3.1 Sigma Points Generation
For state vector $\mathbf{x}\in\mathbb{R}^{n}$ with covariance $\mathbf{P}$, we generate $2n+1$ sigma points:
@@ 3.5 Assocation
[…]
@@
[…]
@@
Remove tracks with time_since_update>max_agetime\_since\_update>max\_age
@@
4 Experiments
4.1 Implementation Details
For all experiments, we employed the same YOLOX [15] detector and the same Re-ID network [17] for SDE-based models, trained with consistent hyperparameter configurations following the established protocols from ByteTrack [52] and BoT-SORT [1]. All models were trained on the MFT25 training set using a NVIDIA A100 GPU, with performance evaluated on the test set using standard MOT metrics, including the comprehensive HOTA [30] metric alongside traditional CLEAR [3] metrics such as MOTA, IDF1, and ID switches (IDs).
4.2 Benchmark Results
@@ 4.3 Ablation Studies
5 Conclusion
In this paper, we introduces a unified underwater MFT benchmark and a specialized tracking framework for fish morphology and erratic swimming patterns. Our baseline achieves state-of-the-art performance with 34.1 HOTA, significantly outperforming other trackers. Statistical analysis reveals fundamental differences between fish and terrestrial tracking scenarios, highlighting the necessity for specialized underwater approaches. However, significant challenges remain in handling visually similar fish appearances, extreme density scenarios, and highly erratic swimming patterns.
Appendix.A Appearance Feature Analysis
To further quantify the challenges posed by underwater fish tracking, we analyze the distribution and separability of identity-specific appearance features extracted from MFT25 in comparison to the widely used pedestrian dataset MOT17. We extract per-track feature embeddings using a pretrained ReID model and visualize their distributions using t-SNE projections and cosine similarity heatmaps.
A complementary analysis is shown in Fig. 7, which depicts pairwise cosine similarities across all track embeddings. In MFT25, the heatmap reveals uniformly high similarity scores across identities, indicating limited discriminative cues for visual association. In contrast, MOT17 exhibits a more diagonal-dominant structure, where self-similarity (along the diagonal) is clearly distinguishable from inter-class similarity, reflecting stronger feature separability. These observations underscore the difficulty of re-identification in underwater settings and motivate the need for enhanced appearance modeling techniques specifically tailored to low-contrast, deformable targets.
Figure 7: Pairwise cosine similarity between track-level appearance features. MFT25 shows uniformly high inter-track similarity, while MOT17 exhibits clearer identity separation.
As illustrated in Fig. 8, the track-level embeddings in MFT25 (left) exhibit significantly higher intra-class overlap and reduced inter-class margins compared to MOT17 (right), where distinct identity clusters are more well-formed and linearly separable. This suggests that fish exhibit substantially less distinctive visual signatures, likely due to their similar body shapes, colorations, and the frequent self-occlusions arising from dense schooling behaviors.
Figure 8: t-SNE projection of track-level appearance embeddings. Identity clusters in MFT25 are less separable compared to those in MOT17, reflecting greater visual ambiguity.
Appendix.B Statistical Analysis
Building upon the comparative statistical analysis of motion direction across diverse tracking datasets presented in the main paper, we conducted additional experiments to quantify the distinctive motion characteristics of fish tracking. Our analysis included pedestrians [11], dancers [42], bees [6], and cells [2]. These experiments aim to differentiate fish tracking from other domains based on their unique motion patterns.
Fig. 9 illustrates the angular velocity distribution across different datasets. Our analysis revealed statistically significant differences in motion patterns between domains. The MOT17 [11] dataset demonstrates highly stable angular velocities with predominantly linear motion patterns characteristic of pedestrian movement. Conversely, the MFT25 dataset exhibits substantially greater angular velocity variations and more erratic directional changes while maintaining high swimming speeds. This unpredictable behavior results from the hydrodynamic properties of underwater motion and environmental factors, which generate inherently nonlinear motion trajectories.
@@ 5 Conclusion
As illustrated in Fig. 10, the MFT25 dataset demonstrates substantial fluctuations in average motion speed, whereas the CTMC [2] dataset exhibits relatively stable speed profiles. These fundamental kinematic differences underscore why general-purpose tracking algorithms frequently underperform in underwater environments and emphasize the critical need for specialized fish tracking methodologies capable of adapting to these distinctive motion characteristics.
Appendix.C Per-Sequence Performance
Table 6 presents a comprehensive evaluation of our SU-T baseline across individual sequences in the MFT25 dataset, revealing substantial heterogeneity in tracking difficulty and performance characteristics. The results demonstrate significant variation across sequences, with HOTA scores ranging from 20.3 to 65.2, highlighting the diverse challenges inherent in underwater fish tracking scenarios.
Clip HOTA↑ IDF1↑ MOTA↑ MOTP↑ DetA↑ DetPr↑ DetRe↑ AssA↑ AssPr↑ AssRe↑ IDs↓ IDFP↓ IDFN↓ LocA↑ Frag↓
BT-002 35.961 58.562 55.923 60.727 45.139 53.968 52.020 28.788 43.979 35.754 35 1917 2095 68.930 128
BT-004 51.244 68.933 93.028 69.256 62.416 69.058 68.071 42.125 59.115 47.578 21 2299 2408 75.671 59
MSK-003 34.188 53.067 50.393 60.078 41.026 52.216 47.920 28.930 44.874 35.650 134 12488 15001 68.854 646
PF-002 65.162 98.650 97.300 69.890 64.927 70.203 70.226 65.399 70.729 70.400 0 82 80 76.067 18
SN-009 27.169 28.799 82.766 70.719 58.500 67.828 65.441 12.692 44.586 15.470 173 29339 30853 76.148 431
SN-011 20.264 26.165 47.423 64.313 40.045 48.522 53.843 10.421 41.176 13.032 161 27146 23575 71.909 522
SN-014 40.046 47.718 84.339 72.997 60.300 70.370 67.023 26.700 61.338 29.128 83 9840 10802 77.814 202
Total 33.351 41.717 68.450 67.350 49.943 59.947 59.243 22.425 49.482 26.206 607 83111 84814 73.911 2006
Optimal Conditions.
The PF-002 sequence emerges as an exemplary case, achieving exceptional performance with 65.2 HOTA, 98.7 IDF1, and remarkably zero identity switches. This outstanding performance can be attributed to several favorable environmental and behavioral factors. The sequence features distinct fish appearances with high inter-individual visual variation, enabling robust feature discrimination. Additionally, the predictable movement patterns and minimal occlusions create ideal conditions for consistent trajectory estimation. The near-perfect IDF1 score of 98.7, combined with the absence of identity switches, demonstrates that under optimal conditions, the SU-T approach can achieve human-level tracking accuracy. The low fragmentation count (18) and minimal false positives (82) and false negatives (80) further corroborate the robustness of the tracking framework in favorable underwater environments. The BT-004 sequence represents another success case, achieving strong results with 51.2 HOTA and 68.9 IDF1. This sequence demonstrates robust tracking capabilities in moderately challenging scenarios where fish maintain relatively consistent swimming behaviors. The balanced DetA (62.4) and AssA (42.1) scores indicate that both detection and association components perform adequately, suggesting that the sequence contains manageable levels of visual complexity and motion predictability. The moderate number of identity switches (21) and reasonable fragmentation (59) demonstrate the method’s resilience to typical underwater tracking challenges.
Environmental Complexity.
The MSK-003 sequence exhibits moderate performance with 34.2 HOTA but reveals specific challenges in identity management, suffering from a high number of identity switches (134 IDs). This pattern suggests the presence of crowded underwater environments where multiple fish exhibit similar appearances or engage in frequent close interactions leading to identity confusion. The substantial IDFP (12,488) and IDFN (15,001) values indicate significant challenges in maintaining consistent detections, likely due to dynamic lighting conditions or dense fish populations. The high fragmentation count (646) suggests frequent track interruptions, possibly caused by temporary occlusions or fish moving in and out of the camera’s field of view.
Challenging Conditions.
The SN-009 and SN-011 sequences represent the most challenging scenarios in the dataset, with HOTA scores of 27.2 and 20.3, respectively. These sequences demonstrate substantial identification failures, with exceptionally high IDFN values of 30,853 and 23,575, revealing fundamental limitations of current tracking approaches in visibility-constrained underwater scenarios. SN-011, in particular, shows the most degraded performance with 20.3 HOTA, attributed to challenging conditions including frequent occlusions and highly erratic fish movements. The extremely low AssA score (10.4) indicates severe difficulties in data association, suggesting that traditional appearance-based and motion-based association strategies fail under these conditions. The high number of false negatives (23,575) implies significant detection failures, potentially due to poor water clarity, extreme lighting variations, or camouflaging effects where fish blend with the background environment. SN-009 exhibits similar challenges with a notably low IDF1 score of 28.8, indicating poor identity preservation throughout the sequence. The substantial difference between DetA (58.5) and AssA (12.7) suggests that while detection performance remains reasonable, the association component fails dramatically, likely due to rapid appearance changes or highly unpredictable movement patterns that violate motion model assumptions.
Detection vs. Association.
A critical insight emerges from examining the relationship between detection accuracy (DetA) and association accuracy (AssA) across sequences. High-performing sequences (PF-002, BT-004) maintain balanced DetA and AssA scores, indicating that both components contribute effectively to overall performance. In contrast, challenging sequences (SN-009, SN-011) show substantial disparities, with DetA significantly outperforming AssA, suggesting that while individual detections remain feasible, temporal coherence and identity maintenance become the primary bottlenecks. The precision-recall balance also varies significantly across sequences. Sequences like PF-002 achieve excellent balance with DetPr (70.2) closely matching DetRe (70.2), while problematic sequences show imbalanced patterns. For instance, SN-011 exhibits higher recall (53.8) than precision (48.5), indicating a tendency toward over-detection that subsequently complicates the association process.
Fragmentation and Temporal Consistency.
The fragmentation analysis reveals important insights about temporal tracking consistency. PF-002’s minimal fragmentation (18) contrasts sharply with MSK-003’s extensive fragmentation (646), highlighting how environmental complexity directly impacts trajectory continuity. High fragmentation correlates strongly with challenging conditions, suggesting that improving track continuation strategies could significantly enhance overall performance in difficult scenarios.
Finally, to visualize the practical effectiveness of SU-T, Fig. 11 presents tracking results across diverse scenarios in the MFT25 dataset. The visualizations demonstrate that SU-T maintains consistent identity assignments despite varying challenges including different fish densities, species morphologies, and environmental conditions.
Appendix.D Hyperparameters
We know that the tuning of hyperparameters has a significant impact on experimental results. To ensure robust detection and tracking in challenging underwater fish tracking conditions, we configured task-specific hyperparameters for both the training and inference pipelines. Notably, these hyperparameters were tuned to achieve the best performance on the MFT25 dataset and were kept consistent across all experiments in the main paper to ensure a fair and unbiased comparison.
Detection Configuration.
The detection model employs a depth multiplier of 1.33 and width multiplier of 1.25 to enhance representation capacity for fine-grained fish features. We utilize an input resolution of $800\times 1440$, which aligns with the wide-format aspect ratio commonly observed in underwater recordings. Multi-scale training is applied with random image sizes ranging from 18 to 32 grid units. The model is trained for 80 epochs using SGD with an initial learning rate of $0.001/64$, scheduled with 1 warm-up epoch and 10 final epochs without augmentation to stabilize training. Each image is normalized using ImageNet statistics, with mosaic and mixup augmentations enabled except during the no-augmentation phase.
Tracking Configuration.
We employ a two-stage tracking scheme that combines IoU-based association with ReID embedding similarity. The primary association utilizes a modified IoU metric tailored for fish morphology, denoted as FishIoU, with a threshold of 0.25. A ByteTrack-style secondary association is enabled, with both stages integrated into a temporal consistency module (TCM) using equal weights of 1.0. Track motion smoothing is achieved using an inertia coefficient of 0.05 to accommodate abrupt directional changes while preserving trajectory stability. To enhance identity preservation, we incorporate a ResNet-101-IBN backbone trained specifically on MFT25. ReID-based long-term correction is triggered when similarity exceeds a threshold of 0.4, with confidence-aware embedding gains set to 1.3 for high-confidence detections and 1.2 otherwise.
Evaluation Protocol.
Evaluation is performed every 5 epochs using COCO-style AP metrics. Detection confidence is thresholded at 0.1, and non-maximum suppression (NMS) is applied with an IoU threshold of 0.7.
Acknowledgments
The paper is supported in part by Beijing Smart Agriculture Innovation Consortium Project (BAIC10-2025). The authors gratefully acknowledge National Innovation Center for Digital Fishery - China Agricultural University, Key Laboratory of Agricultural Informatization Standardization - MARA, P. R. China, Key Laboratory of Smart Farming Technologies for Aquatic Animals and Livestock - MARA, P. R. China, National Innovation Center for Digital Agricultural Products Circulation - MARA, P. R. China, and State Key Laboratory of Efficient Utilization of Agricultural Water Resources - China Agricultural University. Weiran Li gratefully acknowledges financial support from the China Scholarship Council (No. 202406350102).